OpenStudy (landon34):

Which of the following expressions represents the least common multiple of 15, 18, and 25? 2 · 3 · 5 2 · 362 · 5^2 2 · 3^3 · 5^3 2 · 3^2 · 5^3

OpenStudy (welshfella):

first get the prime factors 15 = 3*5 18 = 2*3*3 25 = 5*5 The LCM will be the product of these numbers except that where a number occurs more that once in different factors it is only used once. For example a 5 in factors of 15 matches one in factors of 25 so only one of these are used to get the LCM.

OpenStudy (welshfella):

so there are only 2 5's in the answer - that is 5^2 can you find the rest of the LCM?

OpenStudy (landon34):

yeah, I can.

OpenStudy (welshfella):

i guess the 362 should be 3^2 right?

OpenStudy (landon34):

yes.

OpenStudy (welshfella):

where the t and n come from?

OpenStudy (landon34):

oops, I was on the wrong question. srry.

OpenStudy (welshfella):

there are only numbers in the answer....

OpenStudy (landon34):

hold on, let me rewrite the question. Which of the following expressions represents the least common multiple of 15, 18, and 25? 2 · 3 · 5 2 · 3^2 · 5^2 2 · 3^3 · 5^3 2 · 3^2 · 5^3 there, here is the question, the other one had mistakes.